# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

### v4

@@ 2.1.2. Experiment Module
Table 1. Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
GA MFEA MTEA-AD
CI-HS-T1 9.0426e-01 (6.16e-02) 8.4445e-01 (5.51e-02) + 9.3438e-01 (5.03e-02) =
CI-HS-T2 4.2400e+02 (4.90e+01) 2.8119e+02 (5.09e+01) + 3.1099e+02 (4.11e+01) +
CI-MS-T1 5.4562e+00 (1.52e+00) 5.8307e+00 (8.23e-01) - 4.5085e+00 (2.69e-01) +
CI-MS-T2 4.2474e+02 (6.45e+01) 3.1448e+02 (4.41e+01) + 3.1660e+02 (4.39e+01) +
CI-LS-T1 2.1167e+01 (1.82e-01) 2.0234e+01 (8.35e-02) + 2.1086e+01 (2.82e-01) =
CI-LS-T2 4.1476e+03 (6.34e+02) 4.3638e+03 (4.88e+02) = 4.5958e+03 (6.64e+02) -
PI-HS-T1 4.2186e+02 (5.27e+01) 6.6221e+02 (1.11e+02) - 4.4662e+02 (4.13e+01) -
PI-HS-T2 8.6994e+01 (1.93e+01) 8.9970e+01 (1.67e+01) = 1.2177e+02 (2.88e+01) -
PI-MS-T1 5.4736e+00 (1.42e+00) 4.9276e+00 (4.76e-01) = 4.7264e+00 (3.50e-01) +
PI-MS-T2 2.4152e+04 (1.00e+04) 1.0488e+04 (4.77e+03) + 9.1688e+03 (4.69e+03) +
PI-LS-T1 5.6540e+00 (1.88e+00) 1.9583e+01 (2.71e+00) - 5.3895e+00 (9.11e-01) =
PI-LS-T2 1.4062e+01 (2.48e+00) 1.9371e+01 (3.26e+00) - 5.1076e+00 (1.44e+00) +
NI-HS-T1 2.9452e+04 (1.49e+04) 9.7165e+03 (4.93e+03) + 1.3459e+04 (4.77e+03) +
NI-HS-T2 4.5183e+02 (6.26e+01) 3.7757e+02 (7.84e+01) + 3.8559e+02 (6.26e+01) +
NI-MS-T1 9.0462e-01 (6.19e-02) 9.3064e-01 (5.16e-02) = 9.8876e-01 (4.48e-02) -
NI-MS-T2 3.8662e+01 (3.97e+00) 2.7501e+01 (3.26e+00) + 2.4831e+01 (3.76e+00) +
NI-LS-T1 4.2895e+02 (5.41e+01) 7.4411e+02 (1.30e+02) - 4.6340e+02 (4.97e+01) -
NI-LS-T2 4.3308e+03 (5.19e+02) 4.5452e+03 (4.90e+02) = 4.7449e+03 (6.45e+02) -
+ / - / = Base 8 / 5 / 5 9 / 6 / 3
@@
The experiment module offers functionalities for conducting comprehensive experiments and analyzing results. Users can execute multiple independent runs of various algorithms to tackle multiple MTO problems. Upon completion of the runs, data including experimental settings, objective values, and decision variables are recorded and stored. These data serve as the basis for generating metric results, conducting significance tests, and creating schematic plots.
Metrics functions operating on the data are categorized into single-task and multitask metrics. The selection of single-task metrics presents the results of each task of the MTO problem in tabular format within MToP while choosing to multitask metrics displays the results of each MTO problem. Significance test results are also included in the table alongside the metric results. All table data can be exported in formats such as tex, csv, and xlsx, as demonstrated in Table 1.
Convergence plots are generated based on the selected table data to facilitate the comparison of convergence behaviors among different algorithms, as depicted in Fig. 4 (a). The convergence plot function can be applied to both single- and multi-objective problems. Furthermore, for multi-objective MTO algorithms and problems, the Pareto front can be plotted, as illustrated in Fig. 4 (b).
@@ 2.1.3. Data Process Module
Each unique execution within the experiment module is considered a data object and is saved to a data file named <MTOData.mat>. The settings and intermediate results of the experiment are stored in data properties, as illustrated in Table 2. This saved data can subsequently be reloaded by the experiment module of MToP.
With the data process module, the data can be merged or split, enabling data reuse and customization. The processing of data is categorized into three types: independent runs, algorithms, and problems. If several sets of data with fewer independent runs are available, the data can be merged to create new data with a larger number of independent runs. Similarly, in cases where existing data needs augmentation with new algorithms for comparison, it is unnecessary to re-execute all comparative experiments. Instead, only the new algorithms need to be executed, and the resulting data can be merged with the original data. Likewise, if not all algorithms within the data are required, the data can be split and re-merged accordingly. The data processing for problems follows a similar approach, wherein new problems can be added or deleted as needed.
In practical use, the specific operations for splitting and merging data are as follows:
  • Data Splitting: If an experiment involves running both the MFEA and EMEA algorithms on same problems, the data can be split by selecting it in the <Data Tree> and clicking the <Algorithms Split> button to generate separate datasets for each algorithm. Similarly, data can also be split by independent runs or problems using the respective options.
  • Data Merging: If the MFEA and EMEA algorithms were run on same problems in separate experiments, the data can be merged by selecting both datasets and clicking the <Algorithms Merge> button, provided the settings are identical except for the algorithm. This process can also be used to merge datasets by independent runs or problems.
2.2. File Structure
Figure 5. File structure of MToP. The root directory contains the main script file <mto.m> and four subfolders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>.
In the root directory <MTO/> of MToP, you’ll find a script file named <mto.m> and four folders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>. The structure of these folders is depicted in Fig. 5.
The <Algorithms/> folder contains various categories of algorithms, utility functions, and the algorithm base class <Algorithm.m>. Specific algorithm files such as <MFEA.m>, <MO_MFEA.m>, and <GA.m> are stored within their respective classification folders. The <Algorithms/Utils/> folder provides public utility functions for generation operators, environmental selection approaches, constraint handling techniques, etc.
The <Problems/> folder houses various categories of problems, the base function folder, and the problem base class <Problem.m>. Specific problem files like <WCCI20_MTSO1.m>, <CMT3.m>, and <CEC19_MaTMO6.m> are stored under their corresponding classification folders. Real-world application problem files are located in the <Real-World Application/> folder, while the <Problems/Base/> folder contains public base functions for problems.
The <Metrics/> folder contains all result evaluation metrics, organized into subfolders for single-objective <Metrics/Single-objective/> and multi-objective <Metrics/Multi-objective> optimization. Specific metric files include <Obj.m>, <Obj_MTS.m>, and <IGD.m>, among others.
Lastly, the <GUI/> folder houses all files used by the GUI of MToP. Among them, <MTO_GUI.m> serves as the main file of the GUI, while <MTO_CMD.m> provides functionality for executing experiments via the command line.
@@ 2.3.1. Algorithm
Evaluation() Fitness function evaluation
@@
2 caption={Algorithm implementation example for multi-population EA.},
@@
15% initialize multiple populations for all tasks
16pops = Initialization(Algo, Prob, Individual);
17
18while Algo.notTerminated(Prob, pops) % main loop
@@
21 off(index1) = Generation(pops{t});
@@
23 off(index2) = KnowledgeTransfer(pops)
@@
25 off = Algo.Evaluation(off, Prob, t);
@@
27 pops{t} = Selection(pops{t}, off);
@@
2 caption={Algorithm implementation example for multifactorial EA.},
@@
15% initialize populations for multifactorial evolution
16population=Initialization_MF(Algo, Prob, Individual);
17
18while Algo.notTerminated(Prob, pops) % main loop
@@
23 idx = [off.MFFactor] == t;
24 off(idx) = Algo.Evaluation(off(idx), Prob, t);
@@
27 population = Selection(population, off);
28end
@@
All algorithms within MToP inherit from the algorithm base class <Algorithm.m>. The properties and methods of this base class are detailed in Table 3. This base class encapsulates all functions that interface with the GUI, simplifying the implementation process for algorithms, which only need to focus on the evolutionary workflow itself. The MTEAs implemented in MToP are outlined in Table 4. Additionally, the single-task EAs are exclusively implemented using multi-population methods tailored for MTO problems. Specific algorithms belonging to this category are listed in Table 5.
Listings LABEL:list:algo and LABEL:list:algo2 exemplify the implementation of multi-population and multifactorial algorithms respectively. The algorithm labels provided in the second line serve as identifiers for classification within the GUI. The run() function within each class orchestrates algorithm execution, with <Algo> representing the object itself and <Prob> denoting the problem object to be solved. Population initialization occurs at the outset of the run() function, utilizing either the multi-population or multifactorial method provided by MToP.
In Listing LABEL:list:algo, the variable <pops> is of <cell> data type, housing multiple populations as <population1, population2, population3...>. Each individual within <pops> is an <Individual> object, containing properties such as objective value <Obj>, constraint violation <CV>, and decision variables <Dec>. Contrastingly, Listing LABEL:list:algo2 utilizes a list data type for the <population> variable, directly containing individuals as <individual1:factor1, individual2:factor1, individual3:factor2...>, along with an additional property <MFFactor>.
@@
In the context of knowledge transfer, the multi-population algorithm requires the implementation of the KnowledgeTransfer() function to acquire knowledge from other tasks. Conversely, the multifactorial approach achieves knowledge transfer through random mating within the Generation() function. The Generation() function operates on the decision variables <Dec> of offspring individuals and is tailored to specific algorithms. Meanwhile, the Evaluation() function, embedded within the algorithm base class, automatically engages the problem object <Prob> for offspring fitness evaluation. This entails calculating the objective value <Obj> and constraint violation <CV> for all individuals. Following each loop iteration, the environmental selection function Selection() is invoked to update the new population. While MToP offers universal Selection() functions, specific algorithms also have the option to reimplement this function. Subsequently, the code progresses to the next loop and invokes notTerminated() to document changes. The algorithmic structure in MToP is designed to accommodate all EAs for solving MTO problems.
Table 4. Multitask evolutionary algorithms included in MToP. The table lists the algorithm name, optimization objective, task type, framework, constraint handling, and a brief description. SO and MO denote single-objective and multi-objective optimization, respectively. MT and MaT represent multi-task and many-task, respectively. MF and MP represent multifactorial and multi-population, respectively.
Algorithm Objective Task Framework Constraint Description
MFEA (Gupta et al., 2016) SO MT MF Multifactorial evolution-based GA for MTO
LDA-MFEA (Bali et al., 2017) SO MT MF Linearized domain adaptation in MFEA
MFDE&MFPSO (Feng et al., 2017) SO MT MF Multifactorial evolution-based DE and PSO
G-MFEA (Ding et al., 2019) SO MT MF Generalized MFEA with decision variable translation and shuffling
MFEA-GHS (Liang et al., 2019) SO MT MF Genetic Transform and hyper-rectangle search in MFEA
MFEA-DV (Yin et al., 2019) SO MT MF Enhanced MFEA with cross-task search direction
MFEA-II (Bali et al., 2020) SO MT MF MFEA with online transfer parameter estimation
SREMTO (Zheng et al., 2020) SO MT MF Self-regulated multitask framework for varying relatedness among tasks
TLTLA (Ma et al., 2020) SO MT MF MFEA with two-level inter- and intra-task transfer learning
MFEA-AKT (Zhou et al., 2021b) SO MT MF MFEA with adaptive knowledge transfer via multiple crossover operator
ASCMFDE (Tang et al., 2021a) SO MT MF Intertask transfer in aligned subspace for MFDE
AT-MFEA (Xue et al., 2022) SO MT MF Affine transformation-enhanced domain adaptation for MFEA
MFEA-DGD (Liu et al., 2023) SO MT MF MFEA based on diffusion gradient descent
MFEA-VC (Wang et al., 2024) SO MT MF MFEA with contrastive variational auto-encoder
IMEA (Hashimoto et al., 2018) SO MT MP Multi-population framework for MTO through island model
EMEA (Feng et al., 2019b) SO MT MP Explicit knowledge transfer for MTO via autoencoding
MFMP (Li et al., 2020) SO MT MP Multi-population-based adaptive DE for MTO
DEORA (Li et al., 2022c) SO MT MP Adaptive task selection for competitive MTO
MKTDE (Li et al., 2022b) SO MT MP Meta-knowledge transfer-based multitask DE
MTES (Bai et al., 2022) SO MT MP Multitask OpenAI-ES via gradient-free evolution multitasking
BLKT-DE (Jiang et al., 2023b) SO MT MP Multitask DE with block-level knowledge transfer
MTSRA (Li et al., 2023a) SO MT MP Improved adaptive DE with competitive task selection
MTDE-ADKT (Zhang et al., 2024a) SO MT MP Multitask DE with adaptive dual knowledge transfer
CEDA (Zhang et al., 2024b) SO MT MF/MP Constrained multitasking via co-evolution and domain adaptation
CMO-LKT (Ban et al., 2025) SO MT MP Local knowledge transfer-based EA for constrained MTO
MTEA-AD (Wang et al., 2022a) SO MT/MaT MP MTEA with adaptive knowledge transfer via anomaly detection
MTES-KG (Li et al., 2024c) SO MT/MaT MP Multitask ES with knowledge-guided external sampling
MTEA-HKTS (Zhao et al., 2024) SO MT/MaT MP MTEA for solving the problem of transfer targets
BoKT (Jiang et al., 2023a) SO MaT MP Bi-objective knowledge transfer framework for many-task optimization
TRADE (Wu et al., 2023) SO MaT MP Transferable adaptive parameter DE for many-task optimization
TNG-NES (Li et al., 2024b) SO MaT MP Transfer task-averaged natural gradient for efficient many-task optimization
MTEA-SaO (Li et al., 2023b) SO/MO MT/MaT MP Adaptive solver multitask framework with implicit knowledge transfer
KR-MTEA (Cui et al., 2023) SO/MO MT/MaT MP Adaptive multi-task EA based on knowledge reuse
SBO (Liaw and Ting, 2019) SO/MO MaT MP Symbiosis in biocoenosis framework for many-task optimization
MaTDE (Chen et al., 2020) SO/MO MaT MP Many-task DE with adaptive archive-based knowledge transfer
EMaTO-MKT (Liang et al., 2022c) SO/MO MaT MP Multi-source knowledge transfer via local distribution estimation
MO-MFEA (Gupta et al., 2017) MO MT MF Multiobjective MFEA with non-dominated sort and crowding distance
MO-MFEA-II (Bali et al., 2021) MO MT MF Cognizant Multitasking for parameter estimation in MO-MFEA
EMT-PD (Liang et al., 2022b) MO MT MF Two-stage adaptive knowledge transfer based on population distribution
EMT-GS (Liang et al., 2023) MO MT MF Generative adversarial networks for knowledge transfer
MM-DE (Chen et al., 2018) MO MT MP Fast memetic DE for multiobjective MTO
AMT-NSGA-II (Da et al., 2019) MO MT MP Curbing negative influences online for seamless transfer
EMT-ET (Lin et al., 2021) MO MT MP Effective knowledge transfer approach via non-dominated sort
MOMFEA-SADE (Liang et al., 2022a) MO MT MP Subspace alignment and adaptive Differential Evolution
MTEA-D-DN (Wang et al., 2023) MO MT MP Neighborhood as a bridge for decomposition-based knowledge transfer
MTEA-D-TSD  (Li et al., 2024a) MO MT MP Transfer search directions among decomposed subtasks
MTDE-MKTA (Li and Gong, 2025) MO MT MP Multitask DE with multiple knowledge types and transfer adaptation
RVC-MTEA (Li et al., 2024g) MO MT MP Reference vector contribution based MTEA for competitive multiobjective MTO
MTEA-DCK (Li et al., 2025) MO MT MP Multiobjective MTEA with diversity- and convergen-based knowledge transfer
Table 5. Single-task evolutionary algorithms for multitask optimization included in MToP. The table lists the algorithm name, optimization objective, task type, category, constraint handling, and a brief description. SO and MO denote single-objective and multi-objective optimization, respectively. GA denotes genetic algorithm, DE denotes differential evolution, SI denotes swarm intelligence, ES denotes evolution strategy, CH denotes constraint handling, ND denotes non-dominated sorting, ID denotes indicator-based, DC denotes decomposition, and RV denotes reference vector.
Algorithms Objective Task Category Constraint Description
GA (Zbigniew, 1996) SO ST/MT GA Genetic algorithm with SBX crossover and polynomial mutation
Global-GA (Bull and Liu, 2024) SO ST/MT GA GA with global crossover to replace cooperative coevolution
DE (Storn, 1996) SO ST/MT DE Differential evolution algorithm with DE/rand/1/bin operator
jDE (Brest et al., 2006) SO ST/MT DE Self-adapting control parameters in DE
JADE (Zhang and Sanderson, 2009) SO ST/MT DE Adaptive DE with fast convergence performance
rank-DE (Gong and Cai, 2013) SO ST/MT DE Ranking-based mutation operators for DE
SHADE (Tanabe and Fukunaga, 2013) SO ST/MT DE Success-history based parameter adaptation for DE
L-SHADE (Tanabe and Fukunaga, 2014) SO ST/MT DE Linear population size reduction for SHADE
LSHADE44 (Poláková, 2017) SO ST/MT DE L-SHADE with competing strategies applied to constrained optimization
CAL-SHADE (Zamuda, 2017) SO ST/MT DE Adaptive constraint handling technique for L-SHADE
jSO (Brest et al., 2017) SO ST/MT DE Champion of CEC 2017 numerical optimization competition via improved DE
NL-SHADE-RSP (Stanovov et al., 2021) SO ST/MT DE Champion of CEC 2021 numerical optimization competition via improved DE
PSO (Kennedy and Eberhart, 1995) SO ST/MT SI Particle swarm optimization with global and particle best update
CSO (Cheng and Jin, 2015) SO ST/MT SI Competitive swarm optimizer for large-scale optimization
MPA (Faramarzi et al., 2020a) SO ST/MT SI Marine predators algorithm for numerical optimization
EO (Faramarzi et al., 2020b) SO ST/MT SI Equilibrium optimizer for numerical optimization
AO (Abualigah et al., 2021) SO ST/MT SI Aquila optimizer for numerical optimization
KLDE&KLPSO (Jiang et al., 2023c) SO ST/MT DE/SI Knowledge learning-based DE and PSO for numerical optimization
CMA-ES (Hansen and Ostermeier, 2001) SO ST/MT ES Evolution strategy with derandomized covariance matrix adaptation
IPOP-CMA-ES (Auger and Hansen, 2005) SO ST/MT ES Restart CMA-ES with increasing population size
sep-CMA-ES (Ros and Hansen, 2008) SO ST/MT ES Separable CMA-ES for large-scale optimization
xNES-as (Schaul, 2012) SO ST/MT ES Natural ES with adaptation sampling
R1-NES (Sun et al., 2013) SO ST/MT ES A Linear Time natural ES for non-separable functions
xNES&SNES (Wierstra et al., 2014) SO ST/MT ES ES with adaptive natural gradients to update distribution
OpenAI-ES (Salimans et al., 2017) SO ST/MT ES Parallelized ES with standard normal distribution gradients
DES (Arabas and Jagodziński, 2020) SO ST/MT ES Matrix-free covariance matrix adaptation ES
MTV-DE (Mezura-Montes et al., 2007) SO ST/MT CH Multiple trial vectors in DE for handling constraints
ECHT (Mallipeddi and Suganthan, 2010b) SO ST/MT CH Ensemble of constraint handling techniques
FROFI (Wang et al., 2016) SO ST/MT CH Incorporating objective function information into the feasibility rule
C2oDE (Wang et al., 2019) SO ST/MT CH Composite DE for constrained optimization
CORCO (Wang et al., 2020) SO ST/MT CH Utilizing the correlation between constraints and objective function
DeCODE (Wang et al., 2021) SO ST/MT CH Decomposition-based multiobjective approach for constrained optimization
VMCH (Wu et al., 2022) SO ST/MT CH Voting-mechanism-based ensemble of constraint handling techniques
CEDE-DRL (Hu et al., 2023) SO ST/MT CH Deep reinforcement learning assisted co-evolutionary DE
CCEF-ECHT (Li et al., 2024d) SO ST/MT CH Competitive and cooperative ensemble of constraint handling techniques
SPEA2 (Zitzler et al., 2001) MO ST/MT ND Improving the strength Pareto approach for multiobjective optimization
NSGA-II (Deb et al., 2002) MO ST/MT ND Multiobjective GA with non-dominated sort and crowding distance
NSGA-III (Deb and Jain, 2014) MO ST/MT ND Reference-point-based nondominated sorting for multiobjective optimization
SMS-EMOA (Beume et al., 2007) MO ST/MT ID Hypervolume-based selection for multiobjective optimization
MO-CMA-ES (Igel et al., 2007) MO ST/MT ES Covariance matrix adaptation for multiobjective optimization
MOEA/D (Zhang and Li, 2007) MO ST/MT DC Decomposition-based multiobjective evolutionary algorithm
MOEA/D-DE (Li and Zhang, 2009) MO ST/MT DC MOEA/D for complex Pareto sets in multiobjective optimization
RVEA (Cheng et al., 2016) MO ST/MT RV Reference vector-guided EA for many-objective optimization
AR-MOEA (Tian et al., 2018) MO ST/MT ID Adaptive reference points for MOEA/D in multiobjective optimization
MSEA (Tian et al., 2021a) MO ST/MT ND Multistage EA for better diversity preservation in multiobjective optimization
CCMO (Tian et al., 2021b) MO ST/MT CH Coevolutionary framework for constrained multiobjective optimization
LMOCSO (Tian et al., 2020) MO ST/MT SI Efficient large-scale multiobjective optimization based on CSO
CMOCSO (Ming et al., 2023) MO ST/MT SI Competitive and cooperative swarm optimizer for multiobjective optimization

### v5

@@ 3.1.2. Experiment Module
The Experiment Module provides the core functionalities for conducting comprehensive, large-scale experiments and analyzing the resulting data. It is designed to manage and execute batch runs, allowing users to test multiple algorithms across multiple MTO problems for a specified number of independent repetitions. Upon completion, the module systematically records all experimental settings, objective values, and decision variables. A key architectural feature is its support for MATLAB’s Parallel Computing Toolbox, enabling the parallel execution of independent runs to significantly reduce total experimental time.
The module supports saving the complete, raw experimental dataset (MTOData) to a .mat file, which can be fully reloaded into the platform at any time. Once the data is generated, the module offers a suite of tools for post-processing and analysis. This includes functions to automatically calculate performance metrics, which are categorized to support both single-task and multi-task indicators, and to perform statistical significance tests to compare algorithmic performance. All generated tables and test results can be exported to .xlsx, .csv, or .tex formats. Furthermore, it provides granular export options for specific use cases, such as exporting convergence data in .csv format compatible with IOHanalyzer (Wang2022IOHanalyzer) or saving the optimal solution (single-objective MTO) or set of Pareto sets (multi-objective MTO) to .mat files for external study.
Finally, the module provides essential visualization capabilities for interpreting the experimental outcomes. As shown in Fig. 2, this includes generating convergence plots to simultaneously compare the performance trajectories of multiple algorithms across various problems (Fig. 2 (a)), a function applicable to both single- and multi-objective problems. Furthermore, for multi-objective MTO tasks, it can plot the acquired Pareto fronts (Fig. 2 (b)) to visually assess the quality of the final solution sets. Since all visualizations are generated as standard MATLAB figures, users can leverage MATLAB’s built-in plotting tools for further customization, annotation, and export.
@@ 3.1.3. Data Process Module
Each unique execution within the Experiment Module is saved as a standardized data object to a MTOData.mat file, which is described in detail in Section 3.2.4, this file stores all experimental settings, results, and metadata. This saved data can be fully reloaded by the Experiment Module for post-processing or managed by the Data Process Module.
The core function of the Data Process Module is to manage and manipulate these data objects, enabling data reuse and customization. It provides flexible Merge and Split capabilities. These operations can be performed along three distinct dimensions of the dataset: by independent runs (e.g., combining two 10-run sets into one 20-run set), by algorithms (e.g., adding a new algorithm’s results to an existing dataset), or by problems (e.g., splitting a large benchmark suite into smaller subsets). This functionality is crucial for maintaining organized datasets and efficiently conducting comparative studies, such as when a new algorithm needs to be benchmarked against a set of previously executed experiments.
3.2. Project Structure
The project structure of MToP is organized into three main perspectives: class diagram, sequence workflow, file structure, and experiment data structure. These perspectives collectively illustrate the static architecture, dynamic interactions, and organizational layout of MToP.
3.2.1. Class Diagram
Figure 3. Class diagram of MToP. The main classes include Algorithm, Problem, Individual.
Fig. 3 illustrates the class diagram of MToP, which outlines the static architecture of the platform. The design is centered around three core base classes and functions:
  • Individualclass is a data structure encapsulating a single solution, storing its decision variables (Dec), objective values (Obj), and constraint information (Con, CV).
  • Algorithm class is the abstract base for all optimization algorithms with parameters ($FE$ for current function evaluations, $Gen$ for generations), managing the population of Individual objects (via an aggregation “has a” relationship) and the main evolutionary loop. A concrete example, MFEA, is shown inheriting from it. The Algorithm class also holds a “use a” dependency on the Problem class, as it needs to evaluate individuals.
  • Problem class is an abstract base for optimization tasks, defining the essential evaluate() method and problem parameters ($T$ for task, $N$ for population size, $M$ for objective, $D$ for dimension).
  • Function module groups static utility functions such as metrics (e.g., IGD()) and evolutionary operators (e.g., GA_Crossover()).
3.2.2. Sequence Workflow
Figure 4. Sequence diagram of MToP. The main workflow involves MTO_GUI / MTO_CMD, Algorithm, Problem, and Metric.
Fig. 4 depicts the dynamic interaction workflow between the main components when running an experiment:
  • The process is initiated by a User through either the MTO_GUI or MTO_CMD interface.
  • Then, the interface handles the setup phase, configuring the Algorithm and Problem instances by calling setParameter() and setTasks(). Next, the interface invokes the run() method of the Algorithm. The Algorithm performs its Initialization() and then enters the main optimization loop, conditioned on while notTerminated().
  • Inside the loop, the Algorithm executes its Evaluation() method, which subsequently calls the evaluate(individual) method on the Problem object. The Problem returns the calculated Obj, Con values. Following evaluation, the Algorithm calls EvolvePopulation() to generate the next population.
  • Once the loop terminates, the Algorithm returns the bestResult to the GUI/CMD interface. The interface then passes this result to the Metric module for Calculation(). Finally, the computed metricValue is returned and displayed to the user.
3.2.3. File Structure
Figure 5. File structure of MToP. The root directory contains the main script file mto.m and four subfolders: Algorithms/, Problems/, Metrics/, and GUI/.
The file structure of MToP is shown in Fig. 5. The root directory contains the main script file mto.m and four primary subfolders:
  • Algorithms/: it contains various categories of algorithms, utility functions, and the algorithm base class Algorithm.m. Specific algorithm files such as MFEA.m, MO_MFEA.m, and GA.m are stored within their respective classification folders. The Algorithms/Utils/ folder contains multiple subfolders, including Operator/ and Selection/, which organize specific functional modules. The Operator/ subfolder stores variation operators such as crossover and mutation, while the Selection/ subfolder includes mating and environmental selection approaches. Other subfolders provide additional utility functions such as constraint handling and multi-objective optimization techniques.
  • Problems/: it contains various categories of problems, the base function folder, and the problem base class Problem.m. Specific problem files like WCCI20_MTSO1.m, CMT3.m, and CEC19_MaTMO6.m are stored under their corresponding classification folders. Real-world application problem files are located in the Real-World Application/ folder, while the Problems/Base/ folder contains public base functions for problems.
  • Metrics/: it contains all result evaluation metrics, organized into subfolders for single-objective Metrics/Single-objective/ and multi-objective Metrics/Multi-objective optimization. Specific metric files include Obj.m, Obj_MTS.m, and IGD.m, among others.
  • GUI/: it contains all files used by the GUI of MToP. Among them, MTO_GUI.m serves as the main file of the GUI, while MTO_CMD.m provides functionality for executing experiments via the command line.
3.2.4. Experiment Data Structure
The experiment data structure is the standardized format used by MToP to save the complete output of an experimental batch, regardless of whether it is generated by the MTO_CMD.m command-line runner or the MTO_GUI.m graphical interface. All experimental data is encapsulated within a single MATLAB struct variable named MTOData, which is then saved to a .mat file (e.g., MTOData.mat). This self-contained design ensures portability and simplifies downstream data management and post-processing.
@@ 3.2.4. Experiment Data Structure
The high-level properties stored within the MTOData struct are summarized in Table 2. The core fields are organized as multi-dimensional arrays to map clearly onto the experimental design. Let $P$ be the number of problems, $A$ be the number of algorithms, and $R$ be the number of Reps.
  • Problems: This is a $1\times P$ struct array. Each element MTOData.Problems(i) stores the static metadata for the $i$-th problem, including its Name, number of tasks (T), objective dimensions (M), decision variable dimensions (D), and maximum function evaluations (maxFE).
  • Algorithms: This is a $1\times A$ struct array. Each element MTOData.Algorithms(j) stores the metadata for the $j$-th algorithm, such as its Name and a struct of its specific parameters (Para).
  • RunTimes: This is a $P\times A\times R$ numerical matrix, where the element (i, j, k) stores the wall-clock execution time for running algorithm $j$ on problem $i$ during repetition $k$.
  • Results: This is the $P\times A\times R$ struct array containing the core optimization output. Each struct MTOData.Results(i, j, k) contains the detailed results for a single run, aggregated across all $T$ tasks and all $G$ saved convergence points (where $G$ is the Results Num). Each struct contains:
    • Obj: Stores the objective values. For single-objective problems, this is a $T\times G$ numerical array storing the single best value for each task at each convergence point. For multi-objective problems, this is a $T\times 1$ cell array, where each cell Obj{t} contains a $G\times N\times M_{t}$ array representing the full objective values of $N$ solutions.
    • CV: Stores the constraint violation values. This is typically a $T\times G$ array (for single-objective problems) or a $T\times G\times N$ array (for multi-objective problems) storing the summed constraint violation for the corresponding solutions in Obj.
    • Dec: Stores the decision variables, which are populated only if the Save_Dec flag is set to true. For single-objective problems, this is a $T\times G\times D_{t}$ array storing the decision vector of the best solution. For multi-objective problems, it is a $T\times G\times N\times D_{t}$ array storing the decision vectors for all $N$ solutions.
  • Metrics: This field is not generated by the experiment runners (CMD or GUI) themselves but is designed to be populated later by the Experiment Module when metrics are calculated and saved back to the data file.
This consistent, multi-dimensional struct-based layout is fundamental to the platform, as it allows the Data Process Module to reliably merge and split datasets by manipulating the Problems, Algorithms, or Reps dimensions of this structure.
@@ 3.3. Code Pattern
This subsection details the code patterns employed in MToP for implementing algorithms, problems, and metrics. By adhering to these standardized patterns, developers can easily extend the platform with new methods while ensuring compatibility with the existing architecture and GUI.
@@ 3.3.1. Algorithm
Evaluation() Function evaluation with algorithm state update
@@
2 caption={Algorithm implementation example for multi-population MTEA.},
@@
15 % initialize multiple populations for all tasks
16 pops = Initialization(Algo, Prob, Individual);
17 % main loop
18 while Algo.notTerminated(Prob, pops)
@@
21 offspring(index1) = Generation(pops{t});
@@
23 offspring(index2) = KnowledgeTransfer(pops);
@@
25 offspring = Algo.Evaluation(offspring, Prob, t);
@@
27 pops{t} = Selection(pops{t}, offspring);
@@
2 caption={Algorithm implementation example for multifactorial MTEA.},
@@
15 % initialize populations for multifactorial evolution
16 population = Initialization_MF(Algo, Prob, Individual);
17 % main loop
18 while Algo.notTerminated(Prob, population)
@@
23 idx = [offspring.MFFactor] == t;
24 offspring(idx) = Algo.Evaluation(offspring(idx), Prob, t);
@@
27 population = Selection(population, offspring);
28 end
@@
All algorithms within MToP inherit from the algorithm base class Algorithm.m. The properties and methods of this base class are detailed in Table 3. This base class encapsulates all functions that interface with the GUI, simplifying the implementation process for algorithms, which only need to focus on the evolutionary workflow itself. The MTEAs implemented in MToP are outlined in Table LABEL:tab:algo-mt of the supplementary files. Additionally, the single-task EAs are exclusively implemented using multi-population methods tailored for MTO problems. Specific algorithms belonging to this category are listed in Table LABEL:tab:algo-st.
Listings LABEL:list:algo and LABEL:list:algo2 exemplify the implementation of multi-population and multifactorial algorithms respectively. The algorithm labels provided in the second line serve as identifiers for classification within the GUI. The run() function within each class orchestrates algorithm execution, with Algo representing the object itself and Prob denoting the problem object to be solved. Population initialization occurs at the outset of the run() function, utilizing either the multi-population or multifactorial method provided by MToP. Here, most standard algorithms read the N property from the Prob object (see Table 4) to use the problem’s default population size, which ensures consistency for comparisons. However, this is not a rigid requirement: algorithms known for dynamic population sizing (e.g., L-SHADE, IPOP-CMA-ES) are implemented to override this default and manage their own population sizes as per their original designs.
In Listing LABEL:list:algo, the variable pops is a cell array. It stores multiple populations grouped by task, in the form of {pop1, pop2, pop3, ...}. Each element pop is itself a separate object array containing all individuals [ind1, ind2, ind3, ...] for a specific task. Each Individual object within these lists contains properties such as Obj (objective value), Con (constraint value), and Dec (decision variables). In contrast, Listing LABEL:list:algo2 uses a object array data type for the whole population variable. This adopts a mixed structure, storing individuals from all tasks together in a single list. To differentiate which task each individual belongs to, the Individual objects in this listing include an additional property, MFFactor. Therefore, the structure of this single list can be conceptualized as [ind1, ind2, ind3, ...], where each individual’s MFFactor property is set to its corresponding task ID (e.g., ind1.MFFactor = 1, ind2.MFFactor = 2, ind3.MFFactor = 1, ...).
@@
In the context of knowledge transfer, the multi-population algorithm requires the implementation of the KnowledgeTransfer() function to acquire knowledge from other tasks. Conversely, the multifactorial approach achieves knowledge transfer through random mating within the Generation() function. The Generation() function operates on the decision variables Dec of offspring individuals and is tailored to specific algorithms.
The Evaluation() method serves as a state-updating wrapper in Algo. It performs two sequential operations: first, it invokes the evaluate() method from the Prob object to calculate the actual Obj and Con values. Second, it updates the algorithm’s internal state properties, such as incrementing the FE counter.
@@
Following, the environmental selection function Selection() is invoked to update the new population. While MToP offers universal Selection() functions, specific algorithms also have the option to reimplement this function. Subsequently, the code progresses to the next loop and invokes notTerminated() to document changes. The algorithmic structure in MToP is designed to accommodate all EAs for solving MTO problems.